OpenStudy (anonymous):

Two vectors A and B have precisely equal magnitudes. In order for the magnitude of A+ B to be 110 times larger than the magnitude of A-B , what must be the angle between them?

OpenStudy (anonymous):

178.96 degrees

OpenStudy (anonymous):

 \[\left| A+B \right|=2Acos(\theta)\] \[\left| A-B \right|=2Asin(\theta)\] therefore we require \[\tan(\theta)=\frac{ 1 }{ 110 }\] the angle between A and B is \[180-2\theta\]